iosangel:

A square has an area of 324 𝑖𝑛^2

iosangel:

Tranquility:

Area = side^2 They gave you the area. How would you find the length of the side?

iosangel:

^2

Tranquility:

So we have the length of the side We want to find the area We would square the length of the side (that's what s^2 is) And we would get the area The question wants us to do the OPPOSITE We have the area, but how would we find the length of the side? What is the opposite of squaring something? ^2 is called squaring but essentially when you have something^2 it is multiplication so something^2 = something * something something^3 = something * something * something you just multiply the number by itself for 2 or 3 or whatever the exponent is

iosangel:

Oh so itd be D

Tranquility:

How did you get to that?

Tranquility:

Do you know what squaring is and what square roots are?

iosangel:

Sort of

Tranquility:

What is 4^2

iosangel:

16

iosangel:

4*4

Tranquility:

Good! What is the squareroot of 16? \(\Large \sqrt[2]{16}=?\)

iosangel:

4

Tranquility:

Perfect! So you realize that \(\Large\sqrt[2]{x^2}\) is going to be EQUAL to x So if we squared a number, we would take the square root of it to find out what that original number is

iosangel:

Ok

Tranquility:

Have you figured out the answer now?

iosangel:

B

Tranquility:

Correct!

iosangel:

Ok thank you

Tranquility:

Of course, no problem!